# To Offload or Not To Offload: Model-driven Comparison of Edge-native and On-device Processing

### v2

@@ Abstract
Computational offloading is a promising approach for overcoming resource constraints on client devices by moving some or all of an application’s computations to remote servers. With the advent of specialized hardware accelerators, client devices are now able to perform fast local processing of specific tasks, such as machine learning inference, reducing the need for offloading computations. However, edge servers with accelerators also offer faster processing for offloaded tasks than was previously possible. In this paper, we present an analytic and experimental comparison of on-device processing and edge offloading for a range of accelerator, network, and application workload scenarios, with the goal of understanding when to use local on-device processing and when to offload computations. We present models that leverage analytical queuing results to capture the effects of dynamic factors such as the performance gap between the device and edge server, network variability, server load, and multi-tenancy on the edge server. We experimentally demonstrate the accuracy of our models for a range of hardware and application scenarios and show that our models achieve a mean absolute percentage error of 2.2% compared to observed latencies. We use our models to develop an adaptive resource manager for intelligent offloading and show its efficacy in the presence of variable network conditions and dynamic multi-tenant edge settings.
publicationid: pubid:
I Introduction
@@
Over the past decade, cloud computing has become the predominant approach for running online services in domains ranging from finance to entertainment. In recent years, a new class of online services has emerged that requires low-latency processing to meet end-user requirements. Examples of such applications include autonomous vehicles, interactive augmented and virtual reality (AR/VR), and human-in-the-loop machine learning inference [1]. Edge computing, which is a complementary approach to cloud computing, is a promising method to meet the needs of such latency-sensitive applications. Conventional wisdom has held that since edge resources are deployed at the edge of the network and are closer to end-users, edge-native processing can provide lower end-to-end latencies than cloud processing. However, recent research [2, 3] has shown that this conventional wisdom does not always hold—in certain scenarios, edge processing can yield worse end-to-end latency despite its network latency advantage over the cloud. In particular, since edge clusters are resource-constrained relative to the cloud, research has shown resource bottlenecks can arise in the presence of time-varying workloads and increase the processing latency at the edge, negating its network latency advantage. Further, such bottlenecks can arise even at moderate levels of utilization, requiring careful resource management to maintain the benefits of edge-native processing [3].
@@
A related trend is the emergence of computational offloading from local devices to remote resources such as those at the edge. [^1] Computational offloading to the cloud is also common, but we focus on edge offloading in this paper. Computational offloading [4, 5] involves using nearby edge resources to take on some, or all, of an application’s processing demands. Such offloading is performed to overcome resource constraints on a local device by leveraging more abundant processing capacity at edge servers. Computational offloading is well-established in mobile computing where mobile devices, which are resource- or battery-constrained, offload local computations to edge servers [6]. In this case, local devices pay a so-called mobility penalty [7], which is the additional network latency to access edge servers, but then benefit from faster edge processing that outweighs this latency overhead. Offloading is also common for Internet-of-Things (IoT) devices that have limited on-board processing abilities and rely on remote processing to implement data processing tasks [8]. In these offloading scenarios, conventional wisdom holds that remote edge processing is faster than local processing despite the network overhead of accessing remote resources due to significant resource constraints at local devices.
@@
With the advent of hardware accelerators in recent years, the conventional wisdom about the benefits of computational edge offloading for local devices needs to be rethought. For example, the rise of the so-called AI PC with onboard AI accelerators [9] has enabled local AI processing for many applications without a need for edge offloading. Similarly, mobile phones are equipped with neural accelerators [10] that enable efficient local inference, while IoT devices such as cameras have specialized accelerators to perform visual tasks (e.g., facial recognition) locally [11]. At the same time, edge servers can also be equipped with more powerful accelerators, such as GPUs, enabling offloaded tasks to also run more efficiently or faster than before. Thus, in the era of programmable accelerators, the question of whether to offload computations from local devices to the edge and when it is advantageous to do so needs to be reexamined. In particular, such a reexamination needs to address several questions. (i) Under what scenarios will local on-device processing using accelerators outperform edge offloading? (ii) Are there still scenarios where edge offloading has an advantage over local processing? (iii) What role does a fast (or slow) network between the device and the edge play in terms of whether to offload or not offload?
@@
Motivated by the above questions, this paper takes an analytic model-driven approach to compare local and edge processing in the era of accelerators. Our hypothesis is that there is no a priori clear advantage of edge processing over on-device processing in the presence of accelerators, and that the choice of where to perform the processing depends on hardware, network, and workload characteristics. In developing our analytic models and validating this hypothesis, this paper makes the following contributions.
@@
  • We use queuing theory-based models to model the behavior of device and edge accelerators when processing requests for local on-device processing and edge offloading scenarios. Our modeling follows a two-level approach: it takes estimated service times obtained through profiling or other prediction techniques as input and embeds them into queuing models to obtain end-to-end response time predictions. Using these models, we develop closed-form analytic bounds to determine when one approach outperforms the other under different hardware, network, and workload scenarios. We also develop analytic models for comparing on-device processing to multi-tenant edge servers that serve multiple clients. Although our primary focus is on offloading, we also show how the models can be extended to account for device-edge collaborative (i.e., split) processing.
@@
  • We experimentally validate our analytic models and bounds using a variety of device and edge accelerators, network configurations, and workloads including deep neural networks (DNNs) and recurrent neural networks (RNNs). Our results show that our models are able to predict the performance behavior of on-device processing and edge offloading under various scenarios, achieving a mean absolute percentage error of 2.2%, with 91.5% of predictions falling within $\pm$5of the observed latency and 100% within $\pm$10%.
  • We develop a dynamic resource manager that incorporates our models and demonstrate how they can be leveraged by this resource manager to adaptively switch between on-device and edge processing under time-varying workloads and under dynamic multi-tenancy scenarios. Our results show that it effectively adapts to changing network conditions and fluctuating request rates at edge servers.
II Background
@@
TABLE I: Characteristics of common programmable accelerators.
Accelerator Peak Power (W) Memory Cost Target Workload
Intel Movidius Myriad X VPU 1 – 2.5 2.5 MB of SRAM \$59 Computer vision and CNN inference
Google Edge TPU 2 8 MB of SRAM \$75 ML Inference with TensorFlow Lite
NVIDIA Jetson TX2 15 8 GB of LPDDR4 \$399 AI and GPU workload
NVIDIA Jetson Orin Nano 15 8 GB of LPDDR5 \$499 AI and GPU workload
NVIDIA A2 GPU 60 16 GB of GDDR6 \$1399 AI and GPU workload
II-A Cloud and Edge Offloading
@@
Computational offloading is a well-known approach where client devices that are resource-constrained offload some or all of their application workload to a remote server. In the case of mobile devices such as AR/VR headsets, tablets, or smartphones, offloading can use nearby edge resources to perform latency-sensitive tasks [4, 5]. While such edge offloading involves a network hop to the edge server (also known as the mobility penalty [7]), subsequent processing at the edge is assumed to be much faster since edge resources are much less constrained than the devices. In the case of IoT devices or ML applications, offloading to the cloud is also common where the abundant processing capacity of the cloud platform is utilized for processing the application workload [12]. For example, large language model (LLM) inference uses GPU clusters in the cloud to answer user queries, while IoT devices use cloud storage and processing to perform analytic and data processing tasks.
In general, computational offloading offers benefits when local processing is constrained and the cost of incurring the additional network latency outweighs the benefits of accessing faster processing in the edge or the cloud.
II-B Hardware Accelerators
@@
In recent years, a wide range of programmable accelerators has been developed to accelerate a range of compute tasks, particularly machine learning compute. Such accelerators have been deployed on client devices, edge, and cloud servers, enabling faster processing of accelerated tasks at each tier. Table 1 depicts the characteristics of several commonly deployed device and edge accelerators, ranging from low-end to high-end options. Specifically, the Intel Movidius Myriad X Vision Processing Unit (VPU) is designed to accelerate vision-based applications [13]. Google Edge Tensor Processing Unit (TPU) optimizes ML inference with the TensorFlow Lite framework on the edge [14]. NVIDIA TX2, Orin Nano, and A2 are all tailored for AI workloads such as deep learning and computer vision [15, 16, 17]. For these specific workloads, programmable accelerators can provide substantial performance improvements, delivering server-like capabilities for some compute-intensive tasks even on low-end devices. With their small form factor and energy-efficient design, integrating accelerators into mobile devices has become increasingly common. For example, Apple devices include a Neural accelerator to accelerate machine learning tasks locally on the device [10]. With the presence of accelerators, fast on-device processing of tasks such as machine learning inference has become feasible without the need to resort to edge offloading. Meanwhile, edge servers can also be equipped with GPUs, which serve as general-purpose accelerators for a range of tasks, or special-purpose accelerators for specific tasks. This enables offloaded tasks to also be accelerated at the edge, making edge offloading useful for more compute-intensive tasks that a device may not be able to handle using local resources.
II-C On-device Processing versus Edge Offloading
@@
As noted above, the availability of accelerators on client devices has enabled on-device processing of many tasks without the need to offload those tasks to the edge or to the cloud. Since local processing does not incur any network hops, it will often be faster than remote processing which incurs an additional network latency. However, since edge servers can also be equipped with GPUs and other accelerators, in some compute-intensive scenarios, edge processing may still provide benefits over local processing on device accelerators. Thus, in the era of accelerators, local on-device processing may be faster in some scenarios, while offloading to edge accelerators may still be faster in other cases (after accounting for the network latency).
@@
Even with extensive offline profiling of execution times on both the device and the edge, selecting the optimal execution strategy remains challenging due to dynamic real-world conditions. On the one hand, mobile devices are subject to network variability due to fluctuating wireless signal strength, interference, and congestion in shared communication channels. Since edge offloading requires transmitting input data to an edge server, network fluctuations introduce unpredictability in network delays, affecting its overall performance. Additionally, mobile devices running applications with prolonged operating durations, such as live object detection on captured video, may experience battery constraints. When operating in low-power mode, the system may underclock its processing units to conserve energy, resulting in extended execution times. On the other hand, edge servers are multi-tenant systems shared by multiple devices, which can result in interference across users and higher latencies. Since devices may independently decide when and whether to offload tasks, the edge server’s load can fluctuate dynamically based on the number and nature of tasks being offloaded at any given time. Compared to traditional cloud data centers, edge servers are often more resource-constrained, limiting their ability to elastically scale applications under resource pressure. Workload spikes in multi-tenant scenarios can increase queuing delays, significantly increasing end-to-end execution times. Moreover, edge servers are commonly multiplexed across different applications to maximize edge resources utilization. At the same time, GPUs and other programmable accelerators designed for edge servers often lack isolation mechanisms such as GPU virtualization that are available in cloud-based GPUs [18], leading to performance interference between applications. As a result, execution times can become variable, making it challenging to determine whether to offload under dynamic workloads.
@@
The above observations motivate the primary research question addressed in this paper: Considering the hardware processing capabilities of the device and edge server, along with the workload characteristics and real-world dynamics, under what scenarios does one strategy outperform the other? Specifically, does edge offloading still have a role to play when devices are equipped with on-board accelerators? How do workload dynamics, network dynamics, and edge multi-tenancy impact each strategy and their relative performance with respect to one another? To address this question, the next section develops analytic models to capture the end-to-end execution time for each on-device and edge processing, capturing the key factors that impact application performance.
III Model-driven Performance Comparison
@@
In this section, we develop analytical queuing models for edge offloading and on-device processing, and derive closed-form expressions for their expected latencies to compare the two approaches and identify when one outperforms the other.
III-A Analytic Queuing Models
@@
Figure 1: Modeling request execution using (a) edge offloading and (b) on-device processing.
@@
Our analytical models are derived from queuing theory, a well-established mathematical tool for modeling response times experienced by requests in computing systems. Queuing theory has previously been used to model the performance of edge and cloud applications [2, 19, 20], and will serve as the foundation for analyzing edge offloading and on-device processing. As discussed below, we model the end-to-end latency of a request—the duration of time from when the task enters the first queue until it exits the system—for both edge offloading and on-device processing using queuing theory.
@@
Edge offloading. To model the end-to-end request execution time using edge offloading, we need to model the times spent by the request both at the device and at the edge server. We do so by modeling the device and the edge as two separate queuing systems as shown in Figure 1a. We assume that the request arrives at the device and is then forwarded (“offloaded”) to the edge server over the network. We need to model any network queuing delays and the network transmission time of the request from the device to the server. Upon arrival at the edge server, the request is queued for processing and is subsequently scheduled for execution. Since accelerators employ data parallelism to process requests, we model the accelerator at the edge server as a parallel system that processes each request with a degree of parallelism $k_{edge}$. Once the request processing completes, the result is sent back to the device, which incurs network queuing delay and transmission latency on the reverse path.
@@
On-device processing. Similar to edge offloading, we model on-device request processing as a queuing system shown in Figure 1b. Incoming requests generated by the device first enter a local device queue and are then scheduled onto one of the accelerator cores of the device where they are processed with a degree of parallelism $k_{dev}$. Upon completion of processing, the result is returned to the local application and the request exits the system. This model captures an application where a client device generates tasks, utilizes its accelerator cores for local execution, and sends the result to the application.
@@
In queuing theory, the processing time of a request at each queue that it encounters consists of two key components: queuing delay (i.e., the time spent waiting in the queue) and service time (i.e., the time taken to execute the request on the accelerator). Let $w_{dev}^{proc}$, $w_{edge}^{proc}$, $s_{dev}$, and $s_{edge}$ represent the wait times and request service times at the device and edge, respectively. Edge offloading additionally incurs network latency to transmit the request to the edge and the response back to the device, while on-device processing does not incur any network latency since all processing is local. Let $w_{dev}^{net}$ and $w_{edge}^{net}$ denote the respective network queuing delays on the device and the edge. Also, let $n_{req}$ denote the network delay of sending a request from the device to the edge, and $n_{res}$ be the network delay of sending a result from the edge to the device.
@@
$T_{edge}=w_{dev}^{net}+n_{req}+w_{edge}^{proc}+s_{edge}+w_{edge}^{net}+n_{res}$ (1)
@@
III-B Deriving Service Times on Accelerators
Our analytic models require the service time of a workload as input to compute end-to-end response times. The service time reflects the computational demand on the target accelerator and indicates how quickly a request is processed on a given platform. It can be obtained through empirical profiling or a deep learning model that predicts workload execution times on accelerators. For empirical profiling, many applications already support function-level performance logging or provide profiling hooks [21], which can be used to derive the service time of a request by analyzing the corresponding function’s execution metrics. Alternatively, the workload can be profiled on the target hardware using benchmarking tools [22, 23]. In case of model-based prediction, a neural network can be trained to predict the service time of a model on a given target hardware. For example, prior work [24, 25, 26] estimates DNN inference latency by analyzing factors such as network structure and parameter size. For recurrent architectures such as Gated Recurrent Unit (GRU) networks and Long Short-Term Memory (LSTM) networks, existing latency estimation techniques [27, 28] incorporate factors such as temporal dependencies, hidden state dimensionality, and sequence unrolling behavior to accurately predict inference latency. While any of these prior techniques can be used with our analytic models, our experiments use a simple neural network model to predict the service times.
III-C Deriving and Analyzing End-to-End Latency
In this section we compute the expected latency for edge offloading and on-device processing. Our analysis focuses on applications that require processing results on the device such as AR/VR applications. Note that our models can also be extended to capture the performance of applications that require results on the edge. In the following, we leverage classical queuing theory to fundamentally understand the conditions under which one strategy outperforms the other.
III-C1 Comparing Edge Offload and On-Device Execution
We begin our analysis by examining the conditions under which edge offloading results in higher latency compared to on-device processing. For workloads that benefit from hardware acceleration such as those utilizing GPUs, processing times can be significantly reduced compared to processing on general-purpose CPUs. In the following discussion, we refer to such workloads as accelerator-driven workloads. While our analysis primarily focuses on DNN inference, the results generalize to other AI workloads as will be discussed in Section III-E. Table II summarizes the notation used in our models.
TABLE II: Notations used in analytic models.

### v3

@@ Abstract.
Computational offloading is a promising approach for overcoming resource constraints on client devices by moving some or all of an application’s computations to remote servers. With the advent of specialized hardware accelerators, client devices can now perform fast local processing of specific tasks, such as machine learning inference, reducing the need for offloading computations. However, edge servers with accelerators also offer faster processing for offloaded tasks than was previously possible. In this paper, we present an analytic and experimental comparison of on-device processing and edge offloading for a range of accelerator, network, multi-tenant, and application workload scenarios, with the goal of understanding when to use local on-device processing and when to offload computations. We present models that leverage analytical queuing results to derive explainable closed-form equations for the expected end-to-end latencies of both strategies, which yield precise, quantitative performance crossover predictions that guide adaptive offloading. We experimentally validate our models across a range of scenarios and show that they achieve a mean absolute percentage error of 2.2% compared to observed latencies. We further use our models to develop a resource manager for adaptive offloading and show its effectiveness under variable network conditions and dynamic multi-tenant edge settings.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
@@
Over the past decade, cloud computing has become the predominant approach for running online services in domains ranging from finance to entertainment. In recent years, a new class of online services has emerged that requires low-latency processing to meet end-user requirements. Examples of such applications include autonomous vehicles, interactive augmented and virtual reality (AR/VR), and human-in-the-loop machine learning inference (Mosqueira2022). Edge computing, a complementary approach to cloud computing, is a promising method for meeting the needs of such latency-sensitive applications. Conventional wisdom has held that since edge resources are deployed at the edge of the network and are closer to end-users, edge-native processing can provide lower end-to-end latencies than cloud processing. However, recent research (Ali2021; WangB2024) has shown that this conventional wisdom does not always hold—in certain scenarios, edge processing can yield worse end-to-end latency despite its network latency advantage over the cloud. In particular, since edge clusters are resource-constrained relative to the cloud, research has shown that resource bottlenecks can arise under time-varying workloads and increase the processing latency at the edge, negating its network latency advantage and requiring careful resource management to maintain its benefits (WangB2024).
@@
A related trend is the emergence of computational offloading from local devices to remote resources such as those at the edge. [^1] Computational offloading to the cloud is also common, but we focus on edge offloading in this paper. Computational offloading (Satya2009; Davis2004) involves using nearby edge resources to take on some, or all, of an application’s processing demands. Such offloading is performed to overcome resource constraints on a device by leveraging the more abundant processing capacity at edge servers. Computational offloading is well-established in mobile computing where mobile devices, which are resource- or battery-constrained, offload local computations to edge servers (Ren2019). In this case, local devices pay a so-called mobility penalty (Satyanarayanan2019), which is the additional network latency to access edge servers, but then benefit from faster edge processing that outweighs this latency overhead. Conventional wisdom holds that remote edge processing is faster than local processing despite the network overhead of accessing remote resources due to significant resource constraints at local devices.
@@
With the advent of hardware accelerators in recent years, the conventional wisdom about the benefits of computational edge offloading for local devices needs to be rethought. For example, the rise of the so-called AI PC with onboard AI accelerators (intel_ai_pc) has enabled local AI processing for many applications without the need for edge offloading. Similarly, mobile phones are equipped with neural accelerators (apple_neural_engine) that enable efficient local inference, while IoT devices such as cameras have specialized accelerators to perform visual tasks (e.g., facial recognition) locally (netatmo_camera). At the same time, edge servers can also be equipped with more powerful accelerators, such as GPUs, enabling offloaded tasks to also run more efficiently or faster than before. Thus, in the era of programmable accelerators, the question of whether to offload computations from local devices to the edge and when it is advantageous to do so needs to be reexamined. In particular, such a reexamination needs to address several questions. (i) Under what scenarios will local on-device processing using accelerators outperform edge offloading? (ii) Are there still scenarios where edge offloading has an advantage over local processing? (iii) How does multi-tenancy at edge servers affect the decision to offload or execute on-device?
@@
Motivated by the above questions, this paper takes an analytic model-driven approach to compare local and edge processing in the era of accelerators. Our hypothesis is that there is no a priori clear advantage of edge processing over on-device processing in the presence of accelerators, and that the choice of where to perform the processing depends on hardware, network, and workload characteristics. The novelty of this work lies in deriving closed-form equations for the expected end-to-end latency of each strategy, which are explainable and require no workload-specific training. These equations yield precise, quantitative performance-crossover predictions that can be embedded in OS schedulers and resource managers to dynamically decide whether to run an application component on the device or at the edge. In developing our analytic models and validating our hypothesis, this paper makes the following contributions.
@@
  • We develop queuing theory–based models to model the behavior of device and edge accelerators when processing requests in on-device and edge offloading scenarios. Our modeling adopts a two-level approach: it first uses estimated service times obtained through profiling or prediction techniques as input, and then embeds them into queuing models to predict end-to-end response times. We develop closed-form analytic bounds using these models to determine when one approach outperforms the other under different hardware, network, and workload scenarios. We further extend the analysis to model multi-tenant edge servers serving multiple clients, and demonstrate how the models can be extended to capture device–edge collaborative (i.e., split) processing.
@@
  • We experimentally validate our models and bounds across diverse device and edge accelerators, network configurations, and workloads spanning multiple model architectures, including deep neural network (DNN), recurrent neural network (RNN), and large language model (LLM). The results show that our models accurately predict on-device and offloading performance, achieving a 2.2% mean absolute percentage error, with 91.5% of predictions within $\pm$5% and all within $\pm$10% of the observed latency.
  • We develop a resource manager that leverages our models to adapt between on-device processing and edge offloading in response to real-world dynamics. We present two case studies demonstrating that our models enable effective adaptation to changing network conditions and fluctuating request rates at multi-tenant edge servers.
2. Background
@@
2.1. Computational Offloading
@@
Computational offloading is a well-known approach where client devices that are resource-constrained offload some or all of their application workload to a remote server. In the case of mobile devices such as AR/VR headsets, tablets, or smartphones, offloading can use nearby edge resources to perform latency-sensitive tasks (Satya2009; Davis2004). While such edge offloading involves a network hop to the edge server (also known as the mobility penalty (Satyanarayanan2019)), subsequent processing at the edge is assumed to be much faster since edge resources are much less constrained than those of devices. In general, computational offloading offers benefits when local processing is constrained and the benefits of accessing faster processing at the edge or in the cloud outweigh the cost of additional network latency.
2.2. Hardware Accelerators
@@
In recent years, a wide variety of accelerators have been developed to speed up diverse compute tasks, particularly machine learning workloads. Table 1 depicts the characteristics of several commonly deployed device and edge accelerators. For example, NVIDIA TX2, Orin Nano, and A2 are all tailored for AI workloads such as deep learning and computer vision. With their small form factor and energy-efficient design, accelerators are increasingly integrated into mobile devices. For example, Apple devices include a Neural accelerator to accelerate machine learning tasks locally on the device (apple_neural_engine). With accelerators, fast on-device processing of tasks such as machine learning inference has become feasible without resorting to edge offloading. Meanwhile, edge servers can also be equipped with even more powerful GPUs. This enables offloaded tasks to also be accelerated at the edge, making edge offloading useful for more compute-intensive tasks that a device may not be able to handle using local resources.
Table 1. Characteristics of common hardware accelerators.
Accelerator Power Memory Workload
Google Edge TPU (edgetpu2019) 2 W 8 MB ML Inference
NVIDIA Jetson TX2 (tx22017) 15 W 8 GB GPU compute
NVIDIA Jetson Orin Nano (orin2023) 15 W 8 GB GPU compute
NVIDIA A2 GPU (a22021) 60 W 16 GB GPU compute
2.3. On-device Processing versus Edge Offloading
@@
As noted above, the availability of accelerators on client devices has enabled on-device processing of many tasks without the need to offload those tasks to the edge. Since local processing does not incur any network hops, it will often be faster than remote processing which incurs additional network latency. However, since edge servers can also be equipped with GPUs and other accelerators, in some compute-intensive scenarios, edge processing may still provide benefits over local processing on device accelerators. Thus, in the era of accelerators, local on-device processing may be faster in some scenarios, while offloading to edge accelerators may still be faster in others (after accounting for network latency).
@@
Even with extensive offline profiling, selecting the optimal execution strategy remains challenging due to real-world dynamics. On the one hand, mobile devices are subject to network variability such as fluctuations in wireless signal strength. Since edge offloading involves transmitting input data, network fluctuations can introduce unpredictability in network delays, thereby affecting offloading performance. Additionally, devices running applications with prolonged operating durations such as live object detection may face battery constraints. In low-power mode, processors may be underclocked by the OS to conserve energy, extending on-device execution times. On the other hand, edge servers are multi-tenant systems shared by multiple devices, which can lead to interference among clients and higher latencies. Since devices may independently decide when and whether to offload tasks, the edge server’s load can fluctuate dynamically with the number and type of incoming requests. Unlike traditional cloud data centers, edge servers are often more resource-constrained and cannot elastically scale all applications under high load. As a result, workload spikes in multi-tenant scenarios can increase queuing delays and significantly extend end-to-end execution times. Moreover, edge servers are often multiplexed across multiple applications to maximize resource utilization. GPUs and other accelerators designed for edge servers often lack isolation mechanisms such as GPU virtualization that are available in cloud-based GPUs (MIG_nvidia), leading to performance interference between applications. Consequently, execution times at the edge can vary, making it challenging to decide whether to offload under dynamic workloads.
@@
The above observations motivate the primary research question addressed in this paper: Considering the hardware processing capabilities of the device and edge server, along with the workload characteristics and real-world dynamics, under what scenarios does one strategy outperform the other? Specifically, does edge offloading still have a role to play when devices are equipped with on-board accelerators? How do workload dynamics, network dynamics, and edge multi-tenancy impact each strategy and their relative performance with respect to one another? To address these questions, the next section develops analytic models for the end-to-end execution time of edge offloading and on-device processing, capturing the key factors that impact application performance.
3. Model-driven Performance Comparison
@@
In this section, we develop analytic queuing models for edge offloading and on-device processing, and derive closed-form expressions for their expected latencies to compare the two approaches and identify when one outperforms the other.
3.1. Analytic Queuing Models
@@
Figure 1. Modeling request execution using (a) edge offloading and (b) on-device processing.
@@
Our analytic models are derived from queuing theory, a well-established mathematical tool for modeling response times experienced by requests in computing systems. Queuing theory has previously been used to model the performance of edge and cloud applications (Ali2021; Harchol2013; Urgaonkar2005), and here will serve as the foundation for analyzing the end-to-end latency of requests—from the time a task enters the system until it completes—under both edge offloading and on-device processing.
@@
Edge offloading. To model the end-to-end latency of a request under edge offloading, we need to model the times spent by the request both at the device and at the edge server, which we model as two separate queuing systems as shown in Figure 1a. We assume that a request arrives at the device and is then forwarded (“offloaded”) to the edge over the network, incurring network queuing and transmission delays. Upon arrival at the edge, the request is queued for processing and is subsequently scheduled for execution. Since accelerators can process multiple requests in parallel, we model the edge accelerator as a parallel system with a degree of parallelism $k_{edge}$. After processing, the result is sent back to the device, which incurs network queuing and transmission delays on the reverse path.
@@
On-device processing. Similar to edge offloading, we model on-device request processing as a queuing system shown in Figure 1b. Incoming requests generated by the device first enter a local device queue and are then scheduled onto one of the accelerator cores, where they are processed with a degree of parallelism $k_{dev}$. When processing completes, the result is returned to the local application, and the request exits the system. This model captures an application where a client device generates tasks, uses its accelerator cores for local execution, and sends the results back to the application.
@@
In queuing theory, the processing time of a request at each queue that it encounters consists of two key components: queuing delay (i.e., the time spent waiting in the queue) and service time (i.e., the time taken to execute the request on the accelerator). Let $w_{dev}^{proc}$, $w_{edge}^{proc}$, $s_{dev}$, and $s_{edge}$ represent the wait times and request service times at the device and edge, respectively. Edge offloading additionally incurs network latency as requests are sent to the edge and results are returned to the device. Let $w_{dev}^{net}$ and $w_{edge}^{net}$ denote the respective network queuing delays on the device and the edge. Also, let $n_{req}$ denote the network transmission delay of sending a request from the device to the edge, and $n_{res}$ denote the transmission delay of the response along the reverse path.
@@
(1) $T_{edge}=w_{dev}^{net}+n_{req}+w_{edge}^{proc}+s_{edge}+w_{edge}^{net}+n_{res}$
@@
3.2. Deriving Service Times on Accelerators
Our models require the service time of a workload as input to compute end-to-end response times. The service time reflects the computational demand on the target accelerator and represents the time needed to process a request on that platform. It can be obtained either through empirical profiling or via a predictive deep learning model. For profiling, many applications already provide function-level performance logging or profiling hooks (nvidia2023fastertransformer), which can be used to derive service times by analyzing execution metrics. Alternatively, the workload can be benchmarked directly on the target hardware using tools such as (Reddi2020; nvidiasmi). For model-based prediction, a neural network can be trained to predict the service time of a model on a given hardware. For example, prior work predicts DNN inference latency by analyzing network structure and parameter size (Wang2024; Yao2018; Kang2017), RNN inference latency by examining temporal dependencies and sequence unrolling behavior (Ji2024; Li23), and LLM inference latency by using internal layer embeddings to predict remaining output length (Rana2024). While any of these prior techniques can be used with our analytic models, in our experiments we adopt a simple neural network from (Kang2017) to predict service times.
3.3. Deriving and Analyzing End-to-End Latency
In this section we compute the expected latency for edge offloading and on-device processing. Our analysis focuses on applications that require results on the device, such as AR/VR workloads. While the models assume edge offloading returns results to the device which incurs network delay, they can be generalized to applications where results are consumed on the edge (e.g., IoT sensing) by omitting this network delay. In the following, we leverage classical queuing theory to fundamentally understand the conditions under which one strategy outperforms the other.
3.3.1. Comparing Edge Offload and On-Device Execution
We begin our analysis by examining the conditions under which edge offloading results in higher average latency compared to on-device processing. For workloads that benefit from hardware acceleration such as those utilizing GPUs, processing times can be significantly reduced compared to processing on general-purpose CPUs. In the following discussion, we refer to such workloads as accelerator-driven workloads. While our analysis primarily focuses on DNN inference, a representative component commonly found in AR/VR applications, the results generalize to other AI workloads as discussed in Sec 3.5. Table 2 summarizes the notation used in our models.
Table 2. Notations used in analytic models.